1- Bridgeport Construction Company began work on a $407,000 construction contract in 2020. During 2020, Bridgeport incurred costs of $272,000, billed its customer for $217,000, and collected $177,500. At December 31, 2020, the estimated additional costs to complete the project total $157,020.

Prepare Bridgeport’s journal entry to record profit or loss, if any, using (a) the percentage-of-completion method and (b) the completed-contract method.

2- Waterway, Inc. began work on a $7,071,000 contract in 2020 to construct an office building. Waterway uses the completed-contract method. At December 31, 2020, the balances in certain accounts were Construction in Process $1,877,000, Accounts Receivable $237,000, and Billings on Construction in Process $958,000.

Indicate how these accounts would be reported in Waterway’s December 31, 2020, balance sheet
Answer #1

Answer to Question No. 1 :

(a) Percentage of Completion Method

1. Total Estimated Contract Revenue = $ 407,000 (given)

2. Total Costs incurred in 2020 = $ 272,000 (given)

3. Total Estimated Additional Costs = $ 157,020 (given)

4. Therefore Total Estimated Costs = 2. + 3. = $ 429,020

5. Percentage Complete (Cost to Cost Method) = Total Costs incurred / Total Estimated Costs = 2./3. x 100 = 63.40%

6. Therefore Revenue to be recognized for the year 2020 = Total Estimated Contract Revenue x Percentage Complete

= 1. x 5. = $ 407,000 x 63.40% = $ 258, 038

7. Profit/Loss during the Year = Revenue for the year 2020 - Costs incurred for the year 2020 = 6. - 2.

= $ 258,038 - $ 272,000 = ($ 13,962) [Loss]

8. Journal Entry for Recording the Loss =

Construction Expense A/c ............Dr. $ 272,000

To Construction Revenue A/c Cr. $ 258,038

To Construction in Progress A/c Cr. $ 13,962 (loss)

[Being Contract Revenue, Expense and Loss recognized for the year 2020]

(b) Completed Contract Method

1. No Profit/Loss to be recognized for the Year 2020 as the same is to be recognized in the year the contract is completed.

Answer to Question No. 2 :

Waterway Inc
Balance Sheet as at 31st December 2020
Particulars Amount in $
Current Assets
Construction - in - Process            1,877,000
Less: Billings on Construction - in - Process              (958,000)          919,000
Accounts Receivable          237,000